One of the rectangular components of a force is 20N and it makes an angle of 30° withthe force. The magnitude of other component will be??

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Fcos30^{\circ}=20N
\implies F=\dfrac{40}{\sqrt{3}}N

Hence the other component=Fsin30^{\circ}=\dfrac{20}{\sqrt{3}}N